Find the accumulated amount after 10 yr if $5500 is invested at 3%/year compounded continuously.
Answer by stanbon(57246)

Find the accumulated amount after 10 yr if $5500 is invested at 3%/year compounded continuously.
----------------------------
A(t) = P*e^(rt)
------
A(10) = 5500*e^(0.03*10)
---
A(10) = 5500*e^(0.3)
---
A(10) = $7424.22
====================
Cheers,
Stan H.